I need to find all the different triangles that have sides with whole number lengths and a perimeter of 12.
I need to find the possibilities for length and width for a rectangle with an area of 120 cm2 if the side lengths are whole numbers and which possibility gives the smallest perimeter

Answer by checkley77(12844)

2 BY 5 BY 5=12
3 BY 4 BY 5=12
4 BY 4 BY 4=12
----------------------------
LW=120
10*12 OR 20+24=44 FOR PRERIMETER;
5*24 0R 10+48=58
SQRT120=10.95445
10.95445*10.95445 OR 4*10.95445=43.8178 A SQUARE ALWAYS GIVES THE MINIMAL PERIMETER.